# Counterfactual Token Generation in Large Language Models

### v1

@@ Abstract
“Sure, I am happy to generate a story for you: Captain Lyra stood at the helm of her trusty ship, the Maelstrom’s Fury, gazing out at the endless sea. […] Lyra’s eyes welled up with tears as she realized the bitter truth – she had sacrificed everything for fleeting riches, and lost the love of her crew, her family, and herself.” Although this story, generated by a large language model, is captivating, one may wonder—how would the story have unfolded if the model had chosen “Captain Maeve” as the protagonist instead? We cannot know. State-of-the-art large language models are stateless—they maintain no internal memory or state. Given a prompt, they generate a sequence of tokens as an output using an autoregressive process. As a consequence, they cannot reason about counterfactual alternatives to tokens they have generated in the past. In this work, our goal is to enhance them with this functionality. To this end, we develop a causal model of token generation that builds upon the Gumbel-Max structural causal model. Our model allows any large language model to perform counterfactual token generation at almost no cost in comparison with vanilla token generation, it is embarrassingly simple to implement, and it does not require any fine-tuning nor prompt engineering. We implement our model on Llama 3 8B-instruct and conduct both qualitative and quantitative analyses of counterfactually generated text. We conclude with a demonstrative application of counterfactual token generation for bias detection, unveiling interesting insights about the model of the world constructed by large language models.
@@ 1 Introduction
Currently available LLMs are stateless—they maintain no internal memory or state. Given an input prompt, they generate a sequence of tokens [^1] Tokens are the units that make up sentences and paragraphs. Examples of tokens include (sub-)words, symbols, numbers, and special end-of-sequence tokens. as output using an autoregressive process [12, 13]. At each time step, they first use a neural network to map the prompt and the (partial) sequence of tokens generated so far to a token distribution. Then, they use a sampler to draw the next token at random from the token distribution. [^2] Multiple lines of evidence suggest that, if a LLM is forced to output tokens deterministically, its performance worsens [14]. Finally, they append the next token to the (partial) sequence of tokens, and continue until a special end-of-sequence token is sampled. To understand why this autoregressive process is insufficient to reason counterfactually about alternatives to a previously generated sequence of tokens, we will use an illustrative example.
@@
Figure 1: Illustrative examples of autoregressive token generation. In all panels, plain text indicates the input provided to the LLM and highlighted text indicates the output generated by the model. Each token in the output sequence is highlighted in a different color to represent the (stochastic) state of the sampler. Panel (a) shows an LLM’s output to a user’s prompt using vanilla autoregressive token generation. Panels (b, c) show an LLM’s output to an input comprising a user’s prompt and an unmodified/modified part of the original output from Panel (a) using vanilla autoregressive token generation. Panel (d) shows an LLM’s counterfactual output to an input comprising a user’s prompt and a modified part of the output from Panel (a) using autoregressive token generation augmented with the Gumbel-Max SCM.
@@
Consider that we ask an LLM to share its favorite color, as shown in Figure 1(a). Had the LLM chosen a different color (e.g., purple instead of blue), what would the rest of its output have been? To answer such a counterfactual question, we need to implement two actions: (i) modify the (partial) sequence of tokens fed to the neural network used by the LLM and (ii) compel the sampler used by the LLM to behave exactly as it did in the original generation. Using currently available LLMs, we can readily implement the first action, which can be viewed as a causal intervention [15, 16]. We just need to replace “blue” with “purple” in the (partial) sequence of tokens fed to the neural network. However, we cannot easily implement the second action, because the sampler does not specify how it would have behaved after taking the first action while keeping everything else equal. In fact, note that, if we provide the (modified) partial sequence up to and including the world “blue” (“purple”) as input to the LLM, there is no way to ensure that the LLM will generate an output that matches (the structure of) the original output because the (stochastic) state of the sampler is different, as shown in Figures 1(b) and 1(c). [^3] Note that using the same random seed is not sufficient because the inputs in Figure 1(a) and Figures 1(b) and 1(c) differ in their number of tokens.
@@
Our contributions. Our key idea is to augment the autoregressive process underpinning an LLM, particularly the sampler used in the process, using the Gumbel-Max structural causal model (SCM) [17]. Under this model, the sampler is defined through a causal mechanism which receives as an input the distribution of the next token and a set of Gumbel noise values. Importantly, this causal mechanism specifies how the sampler would have behaved under an intervention on the distribution of the next token and thus allows us to answer counterfactual questions about a previously generated sequence of tokens, as shown in Figure 1(d). Along the way, we also introduce an efficient implementation of the augmented autoregressive process that can generate counterfactual tokens at almost no cost in comparison with vanilla token generation. As a proof of concept, we implement our model on Llama 3 8B-instruct, and we conduct experiments to qualitatively and quantitatively analyze the similarity between an LLM’s original output and the one generated via counterfactual token generation. Additionally, we demonstrate the use of our methodology for bias detection, unveiling interesting insights about the model of the world constructed by large language models. An open-source implementation of our model on Llama 3 8B-instruct is available at https://github.com/Networks-Learning/counterfactual-llms.
@@
Further related work. Our work is most closely related to a line of work on counterfactual text generation [18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. In this line of work, given pairs of factual statements and interventions over these statements, the goal is to generate counterfactual statements that match those made by humans—counterfactual statements that are consistent with the underlying model of the world shared by humans. To this end, existing methods typically fine-tune an LLM using a dataset comprising factual statements, interventions over these statements, and counterfactual statements made by humans. In contrast, in our work, our goal is to generate counterfactual statements that are consistent with the underlying model of the world constructed by a given LLM [28, 29, 30, 31]. In this context, our work also relates to a rapidly increasing number of empirical studies assessing the ability of LLMs to answer questions that require counterfactual reasoning [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Here, the LLMs are typically evaluated using multiple choice questions about a given set of factual and counterfactual statements. However, similarly as in the line of work on counterfactual text generation discussed previously, the counterfactual statements are made by humans.
The Gumbel-max structural causal model has previously been used to enable counterfactual reasoning in Markov decision processes [44], temporal point processes [45], and expert predictions [46]. However, to the best of our knowledge, it has not been previously used to enable counterfactual reasoning in LLMs.
@@ 2 A Causal Model of Token Generation
where $S_{q}$ and $\bm{U}=(U_{i})_{i\in\{1,\ldots,K\}}$ are independent exogenous random variables, with $S_{q}\sim P_{Q}$ and $U_{i}\sim P_{U}$, respectively, $f_{D}$ and $f_{T}$ are given functions, $P_{\varnothing}$ is the point mass distribution on $\varnothing$, and $\texttt{last}(S_{i-1})$ denotes the last token of the sequence $S_{i-1}$. Here, the function $f_{D}$ is defined by the transformer architecture of the LLM and the choice of function $f_{T}$ and distribution $P_{U}$ determines the exact mechanism that the LLM’s sampler uses to (stochastically) select the next token $T_{i}$. Note that, there always exists a pair of $f_{T}$ and $P_{U}$ such that the distribution over tokens $D_{i}$ matches the distribution $P^{\mathcal{M}}(T_{i})$ entailed by $\mathcal{M}$ (see Buesing et al. [48], Lemma 2 for a technical argument). Moreover, note that, in the SCM $\mathcal{M}$, the output sequence $S$ contains the prompt sequence to lighten the notation regarding interventions.
@@ 3 Counterfactual Token Generation Using Gumbel-Max SCMs
where $U_{i,v}\sim\text{Gumbel}(0,1)$ are independently distributed Gumbel variables. Importantly, this class of SCMs has been shown to satisfy a desirable counterfactual stability property that can be intuitively expressed as follows. Assume that, at time step $i$, the augmented autoregressive process sampled token $t_{i}$ given $d_{i}=f_{D}(s_{i-1})$. Then, in a counterfactual scenario where $D_{i}=d^{\prime}$, it is unlikely that, at time step $i$, the augmented autoregressive process would have sampled a token $t^{\prime}$ other than $t_{i}$—the factual one—unless, under the token distribution $d^{\prime}$, the relative chance of generating token $t_{i}$ decreased compared to other tokens. More formally, for any token distribution $d^{\prime}\in\Delta(V)$ with $d^{\prime}\neq d_{i}$ such that
@@ 4 Experiments
In this section, we experiment with an implementation of our model on Llama 3 8B-instruct [50], a popular open-weights large language model. We start by qualitatively analyzing an example of counterfactual story generation. Next, we quantitatively analyze the similarity between factual and counterfactual text. We conclude with an application of counterfactual token generation in detecting model biases towards demographic groups. [^6] All experiments ran on an internal cluster of machines, each equipped with 24 Intel(R) Xeon(R) 3GHz CPU cores, 1024GBs of memory and 2 NVIDIA A100 80GB GPUs.
@@ 4.1 How would the story have unfolded for “Captain Maeve”?
    Interventional token generation: it regenerates the second part of the output using vanilla autoregressive token generation, that is, it samples new noise values $u_{j}$ for the second part of the output, as shown in Figure 1(c).
@@
    Counterfactual token generation: it regenerates the second part of the output using Algorithm 1, that is, it reuses the same noise values $u_{j}$ used in the factual generation for the second part of the output, as shown in Figure 1(d).
@@
Figure 2: Examples of factual, interventional and counterfactual stories. Panel (a) shows a factual story, as given by the LLM. Panels (b) and (c) show variants of the story resulting from interventional and counterfactual token generation, respectively. In panels (b), (c), we give as input to the LLM the original prompt along with the first sentence of the factual output (non-highlighted text), modified by replacing “Lyra” with “Maeve”. Blue (green)-highlighted text indicates the tokens of the output that are identical in the factual story and its interventional (counterfactual) counterpart. Red-highlighted text indicates the differences. In both panels, the temperature parameter is set to $\tau=0.9$.
Figures 2(b)2(c) present two alternative versions of the factual story generated using the methods mentioned above. These stories reveal several interesting insights. The story generated with interventional token generation starts diverging from the factual story after only a few tokens, as the method lacks memory of the noise values $u_{j}$ that resulted in the original output. In contrast, the initial part of the counterfactual output remains identical to the factual output, as expected, due to the counterfactual stability property of the Gumbel-Max SCM and the minor nature of changing the protagonist’s name. Although one may expect this to apply for the rest of the counterfactual output, thinking that the protagonist’s name would be irrelevant to the narrative of this particular story, this is not the case. Perhaps surprisingly, the use of “Maeve” instead of “Lyra” results in a partially different output, illustrating that the LLM’s probability distributions over next tokens are sensitive even to minor changes. In Appendix A, we also observe differences between the factual and counterfactual outputs resulting from other seemingly irrelevant interventions, such as changing the name of the ship, removing the adjective “trusty” or replacing the word “sea” with “blue”.
@@
Figure 3: Comparison between interventional and counterfactual token generation. The panels show the edit distance between the factual token sequence and the sequence generated by interventional and counterfactual token generation using (a) the Gumbel-Max SCM defined in Equation 3 and (b) the top-$p$ Gumbel-Max SCM and (c) the top-$k$ Gumbel-Max SCM discussed at the end of Section 3, against various values of the temperature parameter $\tau$, $p$ and $k$, respectively. In panels (b, c) the temperature parameter is set to $\tau=0.6$. In all three panels, the edit distance is averaged over $4{,}000$ output sequences, resulting from two independent interventions per factual sequence, and shaded areas represent $95\%$ confidence intervals.

### v3

@@ Abstract
“Sure, I am happy to generate a story for you: Captain Lyra stood at the helm of her trusty ship, the Maelstrom’s Fury, gazing out at the endless sea. […] Lyra’s eyes welled up with tears as she realized the bitter truth – she had sacrificed everything for fleeting riches, and lost the love of her crew, her family, and herself.” Although this story, generated by a large language model, is captivating, one may wonder—how would the story have unfolded if the model had chosen “Captain Maeve” as the protagonist instead? We cannot know. State-of-the-art large language models are stateless—they maintain no internal memory or state. Given a prompt, they generate a sequence of tokens as an output using an autoregressive process. As a consequence, they cannot reason about counterfactual alternatives to tokens they have generated in the past. In this work, our goal is to enhance them with this functionality. To this end, we develop a causal model of token generation that builds upon the Gumbel-Max structural causal model. Our model allows any large language model to perform counterfactual token generation at almost no cost in comparison with vanilla token generation, it is embarrassingly simple to implement, and it does not require any fine-tuning nor prompt engineering. We implement our model on Llama 3 8B-Instruct and Ministral-8B-Instruct, and conduct a qualitative and a quantitative analysis of counterfactually generated text. We conclude with a demonstrative application of counterfactual token generation for bias detection, unveiling interesting insights about the model of the world constructed by large language models.
@@ 1 Introduction
Currently available LLMs are stateless—they maintain no internal memory or state. Given an input prompt, they generate a sequence of tokens [^1] Tokens are the units that make up sentences and paragraphs. Examples of tokens include (sub-)words, symbols, numbers, and special end-of-sequence tokens. as output using an autoregressive process [12, 13]. At each time step, they first use a neural network to map the prompt and the (partial) sequence of tokens generated so far to a token distribution. Then, they use a sampler to draw the next token at random from the token distribution. [^2] Multiple lines of evidence suggest that, if an LLM is forced to output tokens deterministically, its performance worsens [14]. Finally, they append the next token to the (partial) sequence of tokens, and continue until a special end-of-sequence token is sampled. To understand why this autoregressive process is insufficient to reason counterfactually about alternatives to a previously generated sequence of tokens, we will use an illustrative example.
@@
Figure 1: Illustrative examples of autoregressive token generation. In all panels, plain text indicates the input provided to the LLM and highlighted text indicates the output generated by the model. Each token in the output sequence is highlighted in a different color to represent the (stochastic) state of the sampler. [^4] In Section 2, we formally define the state of the sampler as an exogenous noise variable in an SCM.  Panel (a) shows an LLM’s output to a user’s prompt using vanilla autoregressive token generation. Panels (b, c) show an LLM’s output to an input comprising a user’s prompt and an unmodified/modified part of the original output from Panel (a) using vanilla autoregressive token generation. Panel (d) shows an LLM’s counterfactual output to an input comprising a user’s prompt and a modified part of the output from Panel (a) using autoregressive token generation augmented with the Gumbel-Max SCM.
@@
Consider that we ask an LLM to share its favorite color, as shown in Figure 1(a). Had the LLM chosen a different color (e.g., purple instead of blue), what would the rest of its output have been? To answer such a counterfactual question, we need to implement two actions: (i) modify the (partial) sequence of tokens fed to the neural network used by the LLM and (ii) compel the sampler used by the LLM to behave exactly as it did in the original generation. Using currently available LLMs, we can readily implement the first action, which can be viewed as a causal intervention [15, 16]. We just need to replace “blue” with “purple” in the (partial) sequence of tokens fed to the neural network. However, we cannot easily implement the second action, because the sampler does not specify how it would have behaved after taking the first action while keeping everything else equal. In fact, note that, if we provide the (modified) partial sequence up to and including the world “blue” (“purple”) as input to the LLM, there is no way to ensure that the LLM will generate an output that matches (the structure of) the original output, as shown in Figures 1(a)1(b) and 1(c). [^5] Note that using the same random seed is not sufficient because the number of tokens of the input in Figure 1(a) and the number of tokens of the inputs in Figures 1(b) and 1(c) differ.
@@
Our contributions. Our key idea is to augment the autoregressive process of token generation underpinning an LLM, particularly the sampler used in the process, using a structural causal model (SCM) [15]. More specifically, we define the sampler through a causal mechanism that receives as input the distribution of the next token and a set of noise values, which determine the sampler’s (stochastic) state. Importantly, the use of a causal mechanism specifies how the sampler would have behaved under an intervention on the distribution of the next token and thus allows us to answer counterfactual questions about a previously generated sequence of tokens, as shown in Figure 1(d). Further, to instantiate our model, we use the Gumbel-Max SCM [17], an SCM shown to satisfy a desirable counterfactual stability property which, in the context of token generation, favors counterfactual output sequences that share similarities with the original sequence. Along the way, we also introduce an efficient implementation of the augmented autoregressive process that can generate counterfactual tokens at almost no cost in comparison with vanilla token generation. As a proof of concept, we implement our model on Llama 3 8B-Instruct and Ministral-8B-Instruct, and we conduct experiments to qualitatively and quantitatively analyze the similarity between an LLM’s original output and the one generated via counterfactual token generation. Additionally, we demonstrate the use of our methodology for bias detection, unveiling interesting insights about the model of the world constructed by large language models. We conclude with a comprehensive discussion of the limitations of our model, including additional avenues for applications. An open-source implementation of our model on Llama 3 8B-Instruct and Ministral-8B-Instruct is available at https://github.com/Networks-Learning/counterfactual-llms.
@@
Further related work. Our work is most closely related to a line of work on counterfactual text generation [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. In this line of work, given pairs of factual statements and interventions over these statements, the goal is to generate counterfactual statements that match those made by humans—counterfactual statements that are consistent with the underlying model of the world shared by humans. To this end, existing methods typically fine-tune an LLM using a dataset comprising factual statements, interventions over these statements, and counterfactual statements made by humans. In contrast, in our work, our goal is to generate counterfactual statements that are consistent with the underlying model of the world constructed by a given LLM [29, 30, 31, 32]. In this context, our work also relates to a rapidly increasing number of empirical studies assessing the ability of LLMs to answer questions that require counterfactual reasoning [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Here, the LLMs are typically evaluated using multiple choice questions about a given set of factual and counterfactual statements. However, similarly as in the line of work on counterfactual text generation discussed previously, the counterfactual statements are made by humans.
The works by Ravfogel et al. [45] and Bynum and Cho [46], which have been undertaken concurrently and independently of our work, also use SCMs to model the autoregressive process underpinning LLMs as we do. The work by Ravfogel et al. [45] is most closely related to ours; they model autoregressive generation using the Gumbel-Max SCM and generate counterfactual strings to visualize and analyze the effects of interventions within (the network of) an LLM. However, in contrast to our work, they do not explicitly implement the sampler of an LLM as an SCM and use the same set of noise values during factual and counterfactual generation, but they sample the noise values using a posterior distribution inferred from the factual output. The work by Bynum and Cho [46] is fundamentally different to ours. In the context of a specific task, they utilize an SCM to describe and quantify the causal relations between variables determined by the semantics of that specific task. In this context, it is also relevant to note that the Gumbel-Max SCM has previously been used to enable counterfactual reasoning in other domains such as Markov decision processes [47], temporal point processes [48], and expert predictions [49].
@@ 2 A Causal Model of Token Generation
where $S_{q}$ and $\bm{U}=(U_{i})_{i\in\{1,\ldots,K\}}$ are independent exogenous random variables, with $S_{q}\sim P_{Q}$ and $U_{i}\sim P_{U}$, respectively, $f_{D}$ and $f_{T}$ are given functions, $P_{\varnothing}$ is the point mass distribution on $\varnothing$, and $\texttt{last}(S_{i-1})$ denotes the last token of the sequence $S_{i-1}$. Here, the function $f_{D}$ is defined by the architecture and parameters of the LLM and the choice of function $f_{T}$ and distribution $P_{U}$ determines the exact mechanism that the LLM’s sampler uses to (stochastically) select the next token $T_{i}$. Note that, there always exists a pair of $f_{T}$ and $P_{U}$ such that the distribution over tokens $D_{i}$ matches the distribution $P^{\mathcal{M}}(T_{i})$ entailed by $\mathcal{M}$ (see Buesing et al. [51], Lemma 2 for a technical argument). Moreover, note that, in the SCM $\mathcal{M}$, the output sequence $S$ contains the prompt sequence to lighten the notation regarding interventions. For an illustration of the SCM and its causal graph, refer to Figure 2.
@@
Figure 2: Causal graph of our proposed SCM $\mathcal{M}$ for token generation. Boxes represent endogenous random variables and circles represent exogenous random (noise) variables. The value of each endogenous variable is given by a function of the values of its ancestors in the causal graph, as defined by Eq. 1. The value of each noise variable $U_{i}$ is sampled independently from a given distribution $P_{U}$, and it determines the stochastic state of the LLM’s sampler during the generation of token $T_{i}$ (refer to Fig. 4).
@@ 3 Counterfactual Token Generation Using Gumbel-Max SCMs
where $U_{i,t}\sim\text{Gumbel}(0,1)$ are independently distributed Gumbel variables that determine the stochastic state of the LLM’s sampler during the generation of token $T_{i}$ (refer to Fig. 4). Importantly, this class of SCMs has been shown to satisfy a desirable counterfactual stability property that can be intuitively expressed as follows. Assume that, at time step $i$, the augmented autoregressive process sampled token $t_{i}$ given $d_{i}=f_{D}(s_{i-1})$. Then, in a counterfactual scenario where $D_{i}=d^{\prime}$, it is unlikely that, at time step $i$, the augmented autoregressive process would have sampled a token $t^{\prime}$ other than $t_{i}$—the factual one—unless, under the token distribution $d^{\prime}$, the relative chance of generating token $t_{i}$ decreased compared to other tokens. More formally, for any token distribution $d^{\prime}\in\Delta(V)$ with $d^{\prime}\neq d_{i}$ such that
@@
To understand the intuition behind counterfactual stability, consider the following example. Assume that the vocabulary $V$ contains 2 tokens “A” and “B”, the (factual) distribution $d_{i}$ assigns values $0.6$ and $0.4$, respectively, and the Gumbel-Max SCM samples token “A”. Consider an intervention that changes $d_{i}$ to a distribution $d^{\prime}$ that assigns values $0.7$ and $0.3$ to “A” and “B”, respectively. Then, the property of counterfactual stability ensures that, during the counterfactual generation, token “A” is also sampled because, in comparison to the factual generation, the relative odds are higher, i.e., $0.7$ to $0.3$ vs. $0.6$ to $0.4$. In a way, counterfactual stability “prioritizes” the token sampled in the factual generation, maintaining consistency between the factual and counterfactual text. For a further discussion of the stability property and alternatives to the Gumbel-Max SCM, refer to Section 5.
@@ 4 Experiments
In this section, we experiment with an implementation of our model on Llama 3 8B-Instruct [53] and Ministral-8B-Instruct [54], two popular open-weights large language models. We start by qualitatively analyzing an example of counterfactual story generation. Next, we quantitatively analyze the similarity between factual and counterfactual text. We conclude with an application of counterfactual token generation in detecting model biases towards demographic groups. [^8] All experiments ran on an cluster of machines, each equipped with 24 Intel(R) Xeon(R) 3GHz CPU cores, 1024GBs of memory and 2 NVIDIA A100 80GB GPUs.
@@ 4.1 How would the story have unfolded for “Captain Maeve”?
    Interventional token generation (see Fig 1(c)): it regenerates the second part of the output using vanilla autoregressive token generation, i.e., it samples new noise values $u_{j}$ for the second part of the output using different states $r_{j}$ for the random number generator from those used in the factual generation.
@@
    Counterfactual token generation (see Fig 1(d)): it regenerates the second part of the output using Algorithm 1, i.e., it reuses the same states $r_{j}$ for the random number generator and, hence, the same noise values $u_{j}$ as the ones used in the factual generation.
@@
Figure 3: Examples of factual, interventional and counterfactual stories. Panel (a) shows a factual story, as given by Llama 3 8B-Instruct. Panels (b) and (c) show variants of the story resulting from interventional and counterfactual token generation, respectively. In panels (b), (c), we give as input to the LLM the original prompt along with the first sentence of the factual output (non-highlighted text), modified by replacing “Lyra” with “Maeve”. Blue (green)-highlighted text indicates the tokens of the output that are identical in the factual story and its interventional (counterfactual) counterpart. Red-highlighted text indicates the differences. In both panels, the temperature parameter is set to $\tau=0.9$.
Figures 3(b)3(c) present two alternative versions of the factual story generated using the methods mentioned above. These stories reveal several interesting insights. The story generated with interventional token generation starts diverging from the factual story after only a few tokens, as the method lacks memory of the noise values $u_{j}$ that resulted in the original output. In contrast, the initial part of the counterfactual output remains identical to the factual output, as expected, due to the counterfactual stability property of the Gumbel-Max SCM and the minor nature of changing the protagonist’s name. Although one may expect this to apply for the rest of the counterfactual output, thinking that the protagonist’s name would be irrelevant to the narrative of this particular story, this is not the case. Perhaps surprisingly, the use of “Maeve” instead of “Lyra” results in a partially different output, illustrating that the LLM’s probability distributions over next tokens are sensitive even to minor changes. In Appendix A, we also observe differences between the factual and counterfactual outputs resulting from other seemingly irrelevant interventions, such as changing the name of the ship, removing the adjective “trusty” or replacing the word “sea” with “blue”.
@@
Figure 4: Comparison between interventional and counterfactual token generation. The panels show the edit distance between the factual token sequence and the sequence generated by interventional and counterfactual token generation using (a) the Gumbel-Max SCM defined in Equation 3, (b) the top-$p$ Gumbel-Max SCM, and (c) the top-$k$ Gumbel-Max SCM discussed at the end of Section 3, against various values of the temperature parameter $\tau$, $p$ and $k$, respectively. In panels (b, c) the temperature parameter is set to $\tau=0.6$. In all three panels, the edit distance is averaged over $4{,}000$ output sequences, resulting from two independent interventions per factual sequence, and shaded areas represent $95\%$ confidence intervals. The icons Refer to caption and Refer to caption indicate results for Llama 3 8B-Instruct and Ministral-8B-Instruct respectively.